OpenStudy (anonymous):

Alex stores fertilizer for his farm in a rectangular box. The length of the box is 6 feet, width is 3 inches, and height is 9 inches. Each bag of fertilizer contains 0.225 cubic foot of fertilizer. How many bags of fertilizer will fill the box completely? [1 foot = 12 inches] Numerical answers expected!

OpenStudy (anonymous):

This is doable, first what is the volume of the box? When you have that you need to be careful as it will be in cubic inches and the bags come in cubic foot. Need to change one to continue

OpenStudy (anonymous):

Volume of rectangular box?

OpenStudy (anonymous):

The volume of the box is 162

OpenStudy (anonymous):

162 what? 162 pony tails?

OpenStudy (anonymous):

lol. No 162 feet?

OpenStudy (anonymous):

Wait a sec. 162 is totally wrong!

OpenStudy (anonymous):

Write down the parameters of the box Height Width Length

OpenStudy (anonymous):

My lesson says to calculate the volume you go length x width x height. Height 9ft Width 3ft Length 6ft

OpenStudy (anonymous):

Height 9ft Width 3ft Length 6ft Are you sure?

OpenStudy (anonymous):

Yes, if you read the statement and the question it says "The length of the box is 6 feet, width is 3 inches, and height is 9 inches."

OpenStudy (anonymous):

Are we reading the same thing?

OpenStudy (anonymous):

"The length of the box is 6 FEET, width is 3 INCHES, and height is 9 INCHES."

OpenStudy (anonymous):

Oh my bad! I got this. Length - 6ft Width - 0.1ft ??? Height - 0.3ft ???

OpenStudy (anonymous):

1 foot =12 inches 1/12 foot=1inch So 3inches=1/12 *3=1/4foot 9inchees=1/12 *9=3/4foot

OpenStudy (anonymous):

So the volume is 6foot*1/4foot*3/4foot=18/16=9/8=1.125 cubic foot

OpenStudy (anonymous):

So I am wrong? If I am doing this correctly (which I'm probably not) Then the width will be 0.25 and the height is 0.75 ?? I am sorry :/ I am really bad at Algebra.

OpenStudy (anonymous):

What is the solution?

OpenStudy (anonymous):

You will multiply all three so... 6ft * 0.25ft * 0.75ft = 1.125

OpenStudy (anonymous):

Then divide that by 0.225.... 1.125 / 0.225 = 5

OpenStudy (anonymous):

@Andras

OpenStudy (anonymous):

Yes that is correct